OpenStudy (anonymous):

Select any two integers between -12 and +12 which will become solutions to a system of two equations. Write two equations that have your two integers as solutions. Show how you built the equations using your integers. Your solution and equations should not be the same as those of other students or the textbook. There are infinite possibilities. 1.Solve your system of equations by the addition/subtraction method. Make sure you show the necessary 5 steps.

myininaya (myininaya):

have you selected your integers? Use fun ones like 1 and 0

OpenStudy (anonymous):

Ok, so start with -2 and 2 because they are nice numbers. So the coordinate is (-2, 2) Now start with the equation y=mx+b to determine a particular solution to the problem. 2=-2m+b Now we choose 2 values for m and b such that the equation is true. We can choose any values such as m=2 and b=6, or m=0 and b=2. So now we have two potential equations, y=2x+6 and y=0x+2 Now multiply the first equation by -1 and you get -y=-2x-6 and y=2 add, and you get 0=-2x-4 So 2x=-4 and x=-2 Now resubstitute x=-2 into one of the equations, y=2(-2)+6 and you get y=2 Which checks out.

OpenStudy (anonymous):

Sorry! I am still confused! This is the 5 steps my book says to follow. Step 1 Write both equations in the form ax + by = c. Step 2Multiply the second equation by 2 in order to make the coefficients of the x terms equal. Step 3 Subtract the second equation from the first equation to eliminate the x variable. Step 4 Solve the equation for y. Step 5 Select one equation and substitute 4 for y and solve for x. and check. So lost!

myininaya (myininaya):

so say we pick integers 0 and 1 so (0,1) is the interestion of two lines step 1: y=mx+1 (we can choose any slope for this line) y=0x+1=1 so we have a line y=1 we can write in ax+by=c form 0*x+1*y=1 so not we need to find another equation that interects at this point (0,1) like x=0 so we have 1*x+0*y=0 so are 2 lines are 0x+1y=1 and 1x+0y=0 step 2: 0x+1y=1 1x+0y=0 we don't really need the any of coefficients to be the same in this case so we have y=1 from the first equation (we have solved for y) now we substite y=0 in the second equation and get x+0(0)=x=0 lol she/he didn't say how hard it had to be so we could make this as easy as possible which we did

myininaya (myininaya):

i stop labeling my steps im sorry lol